Back to the NIKA2 main page

Daily Reports

We are going to call this run "Cryo 48 bis" and NIKA2 run 35. We are using the same ini files as for NIKA2 run 24, hence no change (we add "bis" only to specify that this is another pool). The reason of this choice is that we would like to be able to easily cross-calibrate the NIKA2 run34 with the data of this NIKA2 run 35.

Oct 15th, Tuesday

Andrew arrived at the telescope. Florian and Longji are stuck at the Barcelona airport due to demonstrations. Stefano and Bilal are here since last week, already.

Afternoon shift (Bilal + Andrew + Stefano)

We start the pool in good weather conditions (tau~0.3), and everything is alright with NIKA2 after maintenance. We start with project 036-19 since the beam is still broad from the afternoon conditions and the PI is okay with it. At the end of the shift, we start observing 092-19 (source MOOJ2206)

Night shift (Stefano)


We take over, ok, I take over :-D and continue with the observation of MOOJ2206 (proj. 092-19). I perform 2.66 repetitions. Then tau increases above ~0.35 on avg, with instability spikes and I interrupt. Follows the usual set of calibrators. It's earlier than usual, but this way I can see how the sky evolves. The trend of tau to increase ceases. I go on with project 093-19 (2.66 repetitions of source XLSSC072). Then tau becomes too high and the target too low (3:45 UT). I switch to 028-19 (source B10) that is good for higher opacity and now is also below 80 deg elevation.

Oct 16th, Wednesday

Morning shift (Andrew + Angel)

We take over at 6am with the telescope in focus and pointing well, and we continue to observe the Taurus B10 field for 028-19 until ~7:00 UT. Tau has been stable around ~0.4 for most of this time. We had a bit of trouble finding a suitable focus after sunrise, and then had some trouble communicating with the telescope. We finally resume observations for project 091-19 on MS0735 at ~9:00 UT.

Night shift (Stefano + Bilal until 23 UT)

After learning from Bilal everything I have never dared to ask about EMIR, tonight I enjoy some 1-pixel observations.

Oct 17th, Thursday

Morning shift (Andrew)

EMIR observations continue all morning, with tau varying between ~0.6 and 1.7. Florian and Longji finally arrive at the telescope. Weather conditions around a 225GHz tau of 1 and strongly varying aren't anymore suitable for EMIR@3mm backup projects. (CK)

Night shift (Stefano & Florian and Bilal)

attachment:P1200791_mod_sm.JPG attachment:P1200759_mod_sm.JPG attachment:P1200749_sm.JPG

Also tonight is dedicated to one-pixel joys ;-) This is becoming an EMIR pool.

Oct 18th, Friday

Morning shift (Andrew and Longji)

EMIR observations: tau is ~0.5 and we are struggling to find any suitable targets. By 8:30 UT, we can't find any projects that would like 3mm only, so observations cease for now. At 9:00 UT we resume. [Just a comment: Projects scheduled in the pool backup queue, have all been selected to be suitable for opacities above 0.5. - CK]

Night shift (Stefano, Florian + Bilal until 21 UT)

Also tonight we explore the wonders of 1-pixel spectroscopy (tau~0.3-0.55, varying). We wish it were possible to use NIKA2 during EMIR pools too. It would be a fair trade. [082-19 offers to be used as filler project with NIKA2 for pwv values up to 7mm - CK][That makes tau(225)<=0.3; they also say that in very stable conditions they accept higher values of tau, but unfortunately we're not in stable conditions - SB]

Oct 19th, Saturday

Morning shift (Andrew and Longji)

We were very happy to resume observations with NIKA2 this morning. We observed a number of the PCC2G sources until ~6.30 UT when we could begin observations on the JINGLE project D02-19 (tau~0.35). Shortly after sunrise, we started having the usual problems with the focus, and elliptical beams at 1mm. Eventually we returned to the project, and in total we completed around half of the project (2.5h). These sources are relatively close to the Sun. At ~11:00 UT, the wind became a problem, and the antenna had to stop at ~11:10 UT.

Night shift (Stefano & Florian + Bilal)

Wind and snow

Oct 20th, Sunday

Morning shift (Andrew and Longji) We observed nothing this morning due to the wind and snow. At ~10:00 UT, remote EMIR observations begin on the interstellar comet.

Night shift (Stefano & Florian + Bilal until 21 UT)

Today four planets waved the sun goodbye:

attachment:P1200932_mod_sm.JPG attachment:P1200969_mod2_labels_sm.jpg

We take over and observe project 089-19 (source Abell 2443) until ~22:30 UT.
We switch to project 199-19 and observe ACT-CL-J0223 until 2 a.m. UT.
In between we perform the usual calibrators cycle.
We then switch to 036-19 to observe L1495, but the source is too high (>80 deg). We run extra calibrations to fill the time until L1495 is at ~76 deg and then we resume it. At this point we realise that scan speed for L1495 is 70 arcsec/sec and therefore it must be below 67 deg. Therefore we postpone it to the morning shift.
We switch to 160-16 and observe NGC891. The scans are 40 minutes each (very long). This might cause problems with the frequency drifting of KIDs. It's pretty stable, though: probably there won't be problems.

Oct 21st, Monday

Morning shift (Andrew and Longji)

We take over as observations of NGC891 are underway for 160-16 (Nearby galaxies LP). The weather conditions are very good, with tau extremely stable at ~0.2. We have to trust that the 2mm focus and pointing are well matched at 1mm for a while, because the make_imbfits script is taking > 30 minutes to process the 1mm data from NGC891.

Eventually, we see the 1mm pointing results and the beams look excellent, as expected - life is good! We observe L1495 North and South for Bilal's project, 036-19.

Afternoon shift (Bilal)

After technical time, the afternoon weather becomes very unstable. We are in the fog.

Night shift (Stefano & Florian + Bilal until 23:30 UT)

We observe project 089-19, source Abell 2443 until 00:45 UT, with a break for calibrators in between.
We do one repetition of source XLSSC072 (project 093-19)
We complete the usual calibrators cycle on Uranus
As the opacity (225 GHz) and the stability of the atmosphere have now degraded (tau varying, spiky, 0.3 to >0.5), we spend the last hour of the shift on continuing the monitoring of PCCS2 sources.

During the evening/night, the Az tracking has developed periodic wiggles, more prominent on odd-numbered sub-scans. Sergio will shown this issue to Juan P. tomorrow during maintenance.

Oct 22nd, Tuesday

Morning shift (Andrew and Longji)

We continue observing the PCCS2 for a little while before the weather degrades completely and we cease observations at ~06:00 UT. The telescope is surrounded by thick fog.

DailyReportsNika2Pool151019 (last edited 2019-10-22 07:27:35 by NikaBolometer)